OpenStudy (klimenkov):

Prove for a medal. The sum of the first \(n\) terms of the geometric sequence \(S_n\). \(b_1\) is the first term, \(q\) is the ratio.

OpenStudy (klimenkov):

Please, write a formula for \(S_n\) and it's proof.

OpenStudy (klimenkov):

It is better, you don't use any sourses to help yourself.

OpenStudy (helder_edwin):

if \(b_n=b_1q^{n-1}\). Let \[ \large S_n=b_1+b_2++b_3\dots+b_n=b_1+b_1q+b_1q^2+\dots+b_1q^{n-1} \] then \[ \large qS_n=b_1q+b_1q^2+\dots+b_1q^n \]

OpenStudy (experimentx):